OpenStudy (anonymous):

Hal's trail mix contains peanuts to raisins in a ratio of 8 to 5. If Hal uses 325 g of raisins, how many grams of peanuts must he use?

OpenStudy (anonymous):

I don't how to find answer at all.

OpenStudy (anonymous):

p/r=8/5, so p/325=8/5, use cross products, and p=520 grams

OpenStudy (anonymous):

thank you so much